# Assessment of Spatio-Temporal Predictors in the Presence of Missing and Heterogeneous Data

### v2

@@ 5.3 Determining the presence of correlation (Q1)
The AZ-whiteness test (13) with $\lambda=1/2$ applied to multiplex graph $\boldsymbol{\mathbf{g}}^{*}$ allows us for assessing the overall presence of correlation; here, large values of $|C_{\lambda=1/2}(\boldsymbol{\mathbf{g}}^{*})|$ denote a pronounced correlation in the residuals.
Instead, the values of statistics $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ with $\lambda=0$ and $\lambda=1$ provide first insights about the presence of temporal and spatial correlations, respectively. This follows from Theorem 1, as $C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})$ and $C_{\lambda=1}(\boldsymbol{\mathbf{g}}^{*})$ share the same distribution under the assumption of independent residuals; therefore comparing their values indicates which – temporal or spatial – elements in the processing pipeline show margins for improvement.
For instance, $|C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})|>|C_{\lambda=1}(\boldsymbol{\mathbf{g}}^{*})|$ with $|C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})|\gg 0$ would indicate temporal (auto)correlation to be more problematic than spatial correlation. Should this be the case, the designer has to intervene on the temporal processing pipeline to improve the model, e.g., by changing the width of the input window or the neural prediction architecture; we refer to (Gao and Ribeiro, 2022; Cini et al., 2023a) for discussions about different ways of combining spatial and temporal processing in deep architectures.
@@ 5.4 Node-level analysis (Q2)
evaluated on subgraph $\boldsymbol{\mathbf{s}}_{\phi}$ extracted from multiplex $\boldsymbol{\mathbf{g}}^{*}=(V^{*},E^{*},\boldsymbol{\mathbf{w}},\boldsymbol{\mathbf{r}})$. Differently from Section 5.3, node score $c_{\lambda}(\phi)$ is a correlation score (20), not a statistic (13); this particular choice for the assessment function follows from the discussion on the score comparability of Section 5.1.
@@
$N(v_{t})$ denotes the set of 1-hop neighbors of node $v_{t}$. Edges in $\boldsymbol{\mathbf{g}}^{*}$ linking nodes in $N(V_{\phi}^{*})\setminus V_{\phi}^{*}$ are excluded, as they do not directly impact any node of interest $v\in\phi$. Special cases analyzing single nodes $v$ derive from (21) by setting $\phi=\{v\}$, as shown at the top of Figure 4; with little abuse of notation, denote $c_{\lambda}(v)\doteq c_{\lambda}(\{v\})$.
Relevant considerations follow from inspecting $c_{\lambda}(v)$. For example, observing large values of $|c_{\lambda}(v)|$ when relying on global predictive models, like standard graph neural networks sharing parameters across nodes, can be a warning on the presence of local effects that would be better described by local, node-specific models (Montero-Manso and Hyndman, 2021; Cini et al., 2023b).
@@ 5.5 Temporal analysis (Q3)
Request Q3 can be addressed by considering subgraphs $\{\boldsymbol{\mathbf{s}}_{t}\}_{t}$ that slice multiplex graph $\boldsymbol{\mathbf{g}}^{*}$ along the temporal axis. For each time step $t$, define time score
@@
determined by the subgraph $\boldsymbol{\mathbf{s}}_{t}$ corresponding to the nodes at time steps $t-1$, $t$, and $t+1$. More specifically, $\boldsymbol{\mathbf{s}}_{t}$ is the subgraph of $\boldsymbol{\mathbf{g}}^{*}$ defined by those edges with at least one ending node in
@@ 5.6 Local spatio-temporal analysis
As the number of nodes and time steps increase, scores (21) and (27) become less effective in discovering correlation existing in confined regions of the space-time; this is a consequence of $c_{\lambda}(\phi)$ and $c_{\lambda}(\omega)$ being averages across all nodes and all time steps, respectively. To address this limitation, we present a strategy to identify local correlation patterns across time and space altogether.
@@
as the correlation score evaluated on subgraph $\boldsymbol{\mathbf{s}}_{t,v}$ constructed from the $k$-hop neighbors of $v_{t}$, for some positive $k$; see the bottom part of Figure 4. Correlation score $c_{\lambda}(t,v)$ is local in the space-time and accounts for the correlation around $(t,v)$. In the following experiments, spatio-temporal scores $\{c_{\lambda}(t,v),v\in V,t=1,\dots,T\}$ are constructed by considering $k=4$ hops around each space-time location $(t,v)$.
@@
AZ-test statistics: $\quad C_{\lambda=1}(\boldsymbol{\mathbf{g}}^{*})=16.3$, $\quad C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})=18.9$, $\quad C_{\lambda=1/2}(\boldsymbol{\mathbf{g}}^{*})=24.3$. Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Scores involved in the analysis of residuals on synthetic data. Scores associated with $\lambda=0$, $1/2$, and $1$ are depicted in blue, green, and red colors, respectively. Top left) Temporal scores $c_{\lambda}(t)$; a moving average is applied to improve readability. Center) Node-level scores $c_{\lambda}(v)$ as both line plots and heatmaps on the graph. Bottom left) Local spatio-temporal score $c_{\lambda}(t,v)$ for $\lambda=1/2$; see Figure 6 for $\lambda=0$ and $1$. Red and blue boxes (sets $A$ and $B$) highlight regions with spatial and temporal correlation, respectively. Values of the AZ-test statistics are reported at the top left of the figure.
Figure 6: Synthetic residuals generated according to Section 6 (left). Scores $c_{\lambda=0}(t,v)$ (center) and $c_{\lambda=0}(t,v)$ (right).
AZ-test statistic: $\quad C_{\lambda=1}(\boldsymbol{\mathbf{g}}^{*})=14.6$, $\quad C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})=13.9$, $\quad C_{\lambda=1/2}(\boldsymbol{\mathbf{g}}^{*})=20.2$. Refer to caption
AZ-test statistic: $\quad C_{\lambda=1}(\boldsymbol{\mathbf{g}}^{*})=14.8$, $\quad C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})=15.5$, $\quad C_{\lambda=1/2}(\boldsymbol{\mathbf{g}}^{*})=21.4$. Refer to caption
Figure 8: Scores involved in the AZ-analysis of residuals on synthetic residuals with heterogeneous. Residuals at different space-time locations are generated from different distributions.
6 Empirical validation of the scores behavior
The first experiment considers a synthetic dataset where we artificially induced temporal and spatial correlation to show that the proposed scores $c_{\lambda}(\cdot)$ are indeed able to detect correlation patterns.
@@ 6.1 Synthetic residuals
The residual signal $\boldsymbol{\mathbf{r}}$ is modeled as Gaussian noise exhibiting spatial and temporal correlations within specific regions defined by the intersection of sets $A$ and $B$, respectively. Outside $A\cup B$ residuals are i.i.d..
The data is generated for 60 nodes and 400 time steps. Set $A$ covers time steps in $[200,400)$ and nodes in $[15,45)$, whereas set $B$ time steps in $[100,300)$ and nodes in $[30,60)$. The considered graph topology and sets $A$ and $B$ are depicted in Figure 5. Starting from i.i.d. white noise $\varepsilon_{t,v}\sim\mathcal{N}(0,1)$, correlation is induced as follows. Within region $A\setminus B$ of only spatial correlation, residuals are obtained as
@@
$r_{t,v}=\varepsilon_{t,v}+\frac{\varepsilon_{t-1,u}+\varepsilon_{t+1,u}}{2}.$ (30)
@@
$r_{t,v}=\varepsilon_{t,v}+\frac{\sum_{u\in N(v)}\varepsilon_{t,u}}{|N(v)|}+\frac{\varepsilon_{t-1,u}+\varepsilon_{t+1,u}}{2}.$ (31)
Finally, for each point $(t,v)$ outside $A\cup B$, uncorrelated residuals $r_{t,v}$ equal $\varepsilon_{t,v}$. An example of correlated residuals generated from this process is depicted in the left part of Figure 6.
@@ 6.2 Analysis of residuals
The values of the AZ-whiteness test statistics $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ are substantially larger than 00, correctly identifying the overall presence of both spatial and temporal correlation.
@@
The magnitude of all node-level scores $c_{\lambda=1}(v)$ and temporal scores $c_{\lambda=1}(t)$ (red dot-dashed lines) are consistent with set $A$ containing spatial correlation.
@@
Similarly, $c_{\lambda=0}(v)$ and $c_{\lambda=0}(t)$ (blue dashed lines) are consistent with $B$ associated with temporal correlation.
@@
$c_{\lambda=1/2}(v)$ and $c_{\lambda=1/2}(t)$ (green solid lines) are consistent with the union set $A\cup B$; local scores $c_{\lambda=1/2}(t,v)$ have similar behavior. Moreover, the scores are larger where both types of correlation are present ($A\cap B$), than where there is only either of the two ($A\setminus B$ and $B\setminus A$).
Also the values of the local spatio-temporal scores $c_{\lambda}(t,v)$ with $\lambda=1$ and $\lambda=0$ are consistent with both sets $A$ and $B$; we refer the reader to Figure 6.

### v3

@@ 5.3 Determining the presence of correlation (Q1)
The AZ-whiteness test (13) with $\lambda=1/2$ applied to multiplex graph $\boldsymbol{\mathbf{g}}^{*}$ allows us to assess the overall presence of correlation; here, large values of $|C_{\lambda=1/2}(\boldsymbol{\mathbf{g}}^{*})|$ denote a pronounced correlation in the residuals. We select the value $\lambda=1/2$ as the default, since it weighs the spatial and temporal components equally; see the definition in (13). Instead, the values of the statistic $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ with $\lambda=0$ and $\lambda=1$ provide first insights about the presence of temporal and spatial correlations, respectively. This follows from Theorem 1, as $C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})$ and $C_{\lambda=1}(\boldsymbol{\mathbf{g}}^{*})$ share the same distribution under the assumption of independent residuals; therefore, comparing these values reveals whether temporal or spatial components in the processing pipeline have greater room for improvement.
For instance, $|C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})|>|C_{\lambda=1}(\boldsymbol{\mathbf{g}}^{*})|$ with $|C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})|\gg 0$ would indicate that temporal (auto)correlation is more problematic than spatial correlation. Should this be the case, the designer may intervene on the temporal processing pipeline to improve the model, e.g., by changing the width of the input window or the neural prediction architecture; we refer to (Gao and Ribeiro, 2022; Cini et al., 2025) for discussions about different ways of combining spatial and temporal processing in deep architectures.
@@ 5.4 Node-level analysis (Q2)
evaluated on subgraph $\boldsymbol{\mathbf{s}}_{\phi}$ extracted from multiplex $\boldsymbol{\mathbf{g}}^{*}=(V^{*},E^{*},\boldsymbol{\mathbf{w}},\boldsymbol{\mathbf{r}})$. Different from Section 5.3, node score $c_{\lambda}(\phi)$ is a correlation score (20), not a statistic (13); this particular choice for the assessment function follows from the discussion on the score comparability of Section 5.1.
@@
$N(v_{t})$ denotes the set of 1-hop neighbors of node $v_{t}$. Edges in $\boldsymbol{\mathbf{g}}^{*}$ linking nodes in $N(V_{\phi}^{*})\setminus V_{\phi}^{*}$ are excluded, as they do not directly impact any node of interest $v\in\phi$. The special case of analyzing single nodes $v$ derive from (21) by setting $\phi=\{v\}$, as shown at the central part of Figure 4; with little abuse of notation, denote $c_{\lambda}(v)\doteq c_{\lambda}(\{v\})$.
Relevant considerations follow from inspecting $c_{\lambda}(v)$. For example, observing large values of $|c_{\lambda}(v)|$ when relying on global predictive models, like standard graph neural networks sharing parameters across nodes, can be a warning on the presence of local effects that would be better described by local, node-specific models (Montero-Manso and Hyndman, 2021; Cini et al., 2023a).
@@ 5.5 Temporal analysis (Q3)
Question Q3 can be addressed by considering subgraphs $\{\boldsymbol{\mathbf{s}}_{t}\}_{t}$ that slice multiplex graph $\boldsymbol{\mathbf{g}}^{*}$ along the temporal axis. For each time step $t$, define time score
@@
determined by the subgraph $\boldsymbol{\mathbf{s}}_{t}$ corresponding to the nodes at time steps t1t-1, $t$, and $t+1$. More specifically, $\boldsymbol{\mathbf{s}}_{t}$ is the subgraph of $\boldsymbol{\mathbf{g}}^{*}$ defined by those edges with at least one ending node in
@@ 5.6 Local spatio-temporal analysis
As the number of nodes and time steps increases, scores (21) and (27) become less effective in discovering correlations existing in confined regions of the space-time. This limitation arises because $c_{\lambda}(\phi)$ and $c_{\lambda}(\omega)$ are averages across all nodes and all time steps, respectively. To address this limitation, we present a strategy for identifying local correlation patterns jointly across time and space.
@@
as the correlation score evaluated on subgraph $\boldsymbol{\mathbf{s}}_{t,v}$ constructed from the neighbors of $v_{t}$; see the bottom part of Figure 4. Correlation score $c_{\lambda}(t,v)$ is local in the space-time and accounts for the correlation around $(t,v)$.
@@
5.7 Practical considerations
To conclude the section, we offer further discussion along with remarks and implementation guidance.
5.7.1 Assumptions
The presented AZ-analysis relies on a single main assumption, A2, about the prediction residuals, which ensures that the test statistic $C_{\lambda}$ and the scoring function $c_{\lambda}$ have zero expected value in the absence of residual correlation. Moreover, $c_{\lambda}$ increases as correlation increases, and can be meaningfully compared across subgraphs. Assumption A1, by contrast, is not required in practice, since the goal of the AZ-analysis is precisely to discover when A1 does not hold; it is used in this paper solely to establish the asymptotic behavior of $C_{\lambda}$ and $c_{\lambda}$. Finally, A3 is a technical assumption relating to the graph structure that enables the asymptotic analysis of the distribution of $C_{\lambda}$. In particular, A3 prescribes that the weights are positive – they are expected to encode the strength of the relations – and bounded. Both conditions can be enforced by simple transformations that restrict the weights to a desired range. We also recall that weights are optional and, for the sake of this analysis, setting all of them equal to $1$ is a valid alternative.
We comment that violations of Assumption A2 introduce a bias that invalidates the confidence regions of the AZ-whiteness test statistic and makes the correlation scores less interpretable. However, as anticipated in Section 4, A2 is mild. When residuals $\boldsymbol{\mathbf{r}}_{t,v}$ are scalar values (i.e., $d_{y}=1$), the equality ${\rm sgn}(\bar{\boldsymbol{\mathbf{r}}}\,\boldsymbol{\mathbf{r}}_{t,v})={\rm sgn}(\bar{\boldsymbol{\mathbf{r}}})\,{\rm sgn}(\boldsymbol{\mathbf{r}}_{t,v})$ holds. In this case, the assumption reduces to requiring $\mathbb{E}_{\boldsymbol{\mathbf{r}}_{t,v}}[{\rm sgn}(\boldsymbol{\mathbf{r}}_{t,v})]=0$. This is equivalent to $\mathbb{P}(\boldsymbol{\mathbf{r}}_{t,v}<0)=\mathbb{P}(\boldsymbol{\mathbf{r}}_{t,v}>0)$, that is, the median of $\boldsymbol{\mathbf{r}}_{t,v}$ is zero. If $m\doteq\text{median}(\boldsymbol{\mathbf{r}}_{t,v})\neq 0$, offsetting the residuals by m-m resolves the issue. In particular, note that $\boldsymbol{\mathbf{r}}$ and $\boldsymbol{\mathbf{r}}-m$ have the same correlation. Moreover, this offset is applied only to assess the presence of correlation and not to measure the prediction accuracy. Conversely, in the multivariate case, Assumption A2 cannot always be satisfied simply by re-centering the residuals. In such cases, we suggest computing correlation scores $c_{\lambda}(\cdot)$ (or statistics $C_{\lambda}(\cdot)$) separately for each of the $d_{y}$ components. This allows enforcing A2 component-wise. The resulting $d_{y}$ values can then be re-aggregated, e.g., by averaging.
5.7.2 Source of correlation and model improvement
While the AZ-analysis highlights where correlation is present in the prediction residuals, it does not by itself identify the underlying cause. In practice, correlation may arise from multiple factors, including non-stationarities in the data-generating process, issues in the acquisition or preprocessing pipeline, or patterns that the predictive model fails to capture. Sections 5.35.6 illustrate examples of how different correlation patterns can be interpreted and addressed. The AZ-analysis should be viewed as a diagnostic tool that guides the practitioner toward aspects of the system that warrant further investigation. Depending on the identified pattern, interventions to improve model quality may involve redesigning the temporal or spatial processing pipelines, revising the training procedure, or adapting how missing and heterogeneous data are handled.
5.7.3 Choice of the parameters
Setting parameter $\lambda$ to either 0 or $1$ allows focusing the analysis on temporal or spatial correlation in isolation, respectively. The AZ-analysis mainly inspects the scores for $\lambda=0$ and $\lambda=1$, which are complementary, as shown throughout the empirical Sections 67, and 8. Varying $\lambda$ within $(0,1)$ effectively provides an interpolation of the scores $c_{0}(\cdot)$ and $c_{1}(\cdot)$ (and similarly for statistics $C_{0}(\cdot)$ and $C_{1}(\cdot)$); this might be convenient for visualization purposes, e.g., when reporting local scores $c_{\lambda}(t,v)$ as in Figures 6 and 13. In particular, $\lambda=1/2$ provides an equal weighting of the temporal component $\widetilde{C}_{{\rm tm}}$ and the spatial one $\widetilde{C}_{{\rm sp}}$ (see (13) and (20)). Unless there is a specific motivation to consider other values, we suggest focusing on $\lambda=0,1/2,1$.
A second parameter of the analysis is the threshold $\gamma$, which is used to identify statistically significant correlations. According to Theorem 1, $\gamma$ can be conveniently derived from the quantiles of the standard Gaussian distribution, regardless of the number of edges (and associated weights) of the input graph; therefore, the same threshold can be used for any graph. However, we suggest using the AZ-whiteness test mainly in preliminary phases to assess the overall presence of correlation (Section 5.3), as scanning and performing multiple hypothesis tests for several subgraphs can lead to an overall probability of type-I errors for the comprehensive analysis that differs significantly from the nominal level of each individual test. In practice, lower thresholds increase the number of detections, while larger thresholds reduce false alarms at the cost of lower sensitivity.
5.7.4 Handling small subgraphs
In sparse graphs, local scores $\{c_{\lambda}(t,v)\}$ based on 1-hop neighborhoods may lead to high score variance. In fact, this phenomenon results from considering subgraphs with very few edges, and can already be observed in Figure 3(b), where the standard deviation decreases as the number of edges increases. In such cases, the subgraph $\boldsymbol{\mathbf{s}}_{t,v}$ can be constructed from the $k$-hop neighborhood for some $k\geq 2$. This ensures that the number of edges grows approximately as $\delta^{k}$, where $\delta$ is the average node degree. For instance, in the experiments of this paper, we considered $k=4$ hops around each space-time location $(t,v)$. The effect of considering different values of $k$ is shown in Figure 7 and discussed in Section 6 below.
5.7.5 Computational and memory complexity
The computational and memory complexities of the AZ-whiteness test statistic $C_{\lambda}(\boldsymbol{\mathbf{s}})$ and the correlation scores $c_{\lambda}(\boldsymbol{\mathbf{s}})$ scale linearly in the number of spatio-temporal edges and in the residual dimension, respectively. Computing either $c_{\lambda}(\boldsymbol{\mathbf{s}})$ or $C_{\lambda}(\boldsymbol{\mathbf{s}})$ for a generic graph $\boldsymbol{\mathbf{s}}$ and $\lambda\in(0,1)$ has complexity linear in the number of edges $|E_{\boldsymbol{\mathbf{s}}}|$. The complexity of computing ${\rm sgn}(\boldsymbol{\mathbf{r}}_{t,u}^{\top}\boldsymbol{\mathbf{r}}_{\tau,v})$ is $O(d_{y})$, and this operation is carried out for each of the $|E_{\boldsymbol{\mathbf{s}}}|$ edges. Since the denominators in (13) and (20) are also computed in $O(|E_{\boldsymbol{\mathbf{s}}}|)$ time, the total complexity is $O(d_{y}|E_{\boldsymbol{\mathbf{s}}}|)$. In a typical spatio-temporal setting with $N$ sensors, $T$ time steps, and average spatial degree $\delta$ (per $\boldsymbol{\mathbf{g}}_{t}$), we have approximately $\delta NT$ spatial edges and N(T1)N(T-1) temporal edges, so $|E_{\boldsymbol{\mathbf{s}}}|=O\bigl((\delta+1)NT\bigr)$ and hence a computational complexity of $O\bigl((\delta+1)NTd_{y}\bigr).$ When $d_{y}$ and $\delta$ are fixed and moderate, the cost is effectively linear in the number of space-time nodes $|V_{\boldsymbol{\mathbf{s}}}|\approx NT$. More specifically, for the correlation scores, the computation is $O(\delta T)$ for the node scores $c_{\lambda}(v)$, $O(\delta N)$ for the time scores $c_{\lambda}(t)$, and $O(\delta^{k})$ for the local scores $c_{\lambda}(v,t)$ computed over $k$ hops. Regarding memory, the test requires access to the residuals and the edges. Storing all residuals costs $O(|V_{\boldsymbol{\mathbf{s}}}|\,d_{y})=O(NTd_{y})$ memory. The edges require $O(|E_{\boldsymbol{\mathbf{s}}}|)$ memory, which is approximately $\delta NT$ but can be reduced to $O(\delta N)$ when the spatial edges do not vary with time. Finally, we note that static graphs enable the use of sparse operations and the parallelization available in modern (graph) deep learning libraries.
6 Empirical validation of score behavior
The first set of experiments considers a synthetic dataset where we artificially induce temporal and spatial correlation to validate the proposed scores $c_{\lambda}(\cdot)$ and show that they are indeed able to detect correlation patterns. After presenting the generation of the synthetic residuals in Section 6.1, the remainder of the section focuses on the AZ-analysis, commenting on each step of the proposed procedure, validating its effectiveness under missing and heterogeneous data, and comparing it with established approaches from the literature. Two real-world use cases follow in Sections 7 and 8.
@@ 6.1 Synthetic residuals
The residual signal $\boldsymbol{\mathbf{r}}$ is modeled as Gaussian noise exhibiting spatial and temporal correlations within specific regions defined by sets $A$ and $B$, respectively. Outside $A\cup B$ residuals are i.i.d..
The data are generated for 60 nodes and 400 time steps. Set $A$ covers time steps in $[200,400)$ and nodes in $[15,45)$, whereas set $B$ covers time steps in $[100,300)$ and nodes in $[30,60)$. The considered graph topology and sets $A$ and $B$ are depicted in Figure 6. Starting from i.i.d. white noise $\varepsilon_{t,v}\sim\mathcal{N}(0,1)$, correlation is induced as follows. Within region $A\setminus B$, where only spatial correlation is present, residuals are obtained as
@@
$r_{t,v}=\varepsilon_{t,v}+\frac{\varepsilon_{t-1,v}+\varepsilon_{t+1,v}}{2}.$ (30)
@@
$r_{t,v}=\varepsilon_{t,v}+\frac{\sum_{u\in N(v)}\varepsilon_{t,u}}{|N(v)|}+\frac{\varepsilon_{t-1,v}+\varepsilon_{t+1,v}}{2}.$ (31)
Finally, for each point $(t,v)$ outside $A\cup B$, uncorrelated residuals $r_{t,v}$ equal $\varepsilon_{t,v}$. An example of correlated residuals generated from this process is depicted in Figure 5.
Figure 5: Synthetic residuals generated as described in Section 6.1.
AZ-whiteness test statistics: $\qquad C_{0}(\boldsymbol{\mathbf{g}}^{*})=16.2$, $\qquad C_{1/2}(\boldsymbol{\mathbf{g}}^{*})=22.9$, $\qquad C_{1}(\boldsymbol{\mathbf{g}}^{*})=16.1$. Refer to caption
Figure 6: Scores involved in the AZ-analysis of residuals on synthetic data. Scores associated with $\lambda=0$, $1/2$, and $1$ are depicted in blue, green, and red colors, respectively. Top-left) Time scores $c_{\lambda}(t)$; a moving average is applied to improve readability. Center) Node-level scores $c_{\lambda}(v)$ as both line plots and heatmaps on the graph. Bottom-left) Local spatio-temporal score $c_{\lambda}(t,v)$ for $\lambda=1/2$; see Figure 7 for $\lambda=0$ and $1$. Red and blue boxes (sets AA and $B$) highlight regions with spatial and temporal correlation, respectively. Values of the AZ-whiteness test statistics are reported at the top left of the figure.
Figure 7: Local scores $c_{\lambda}(t,v)$ computed over different $k$-hop neighborhoods, $k=1,2,4,$ and $6$, on the synthetic data described in Section 6.1. $k=4$ is the value used in the other figures, unless stated otherwise. Unlike the other figures in this section, the aspect ratio is set close to 1 to better visualize the smoothing effects along both axes. The colormap is consistent across the plots.
@@ 6.2 Analysis of residuals
The values of the AZ-whiteness test statistics $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ are substantially larger [^1] For reference, setting $\alpha=10^{-5}$ in (12) leads to $\gamma\approx 4.3$. than 0, correctly identifying the overall presence of both spatial and temporal correlation.
@@
The magnitude of all node-level scores $c_{\lambda=1}(v)$ and time scores $c_{\lambda=1}(t)$ (red dot-dashed lines) is consistent with set $A$ containing spatial correlation.
@@
Similarly, $c_{\lambda=0}(v)$ and $c_{\lambda=0}(t)$ (blue dashed lines) are consistent with $B$ being associated with temporal correlation.
@@
$c_{\lambda=1/2}(v)$ and $c_{\lambda=1/2}(t)$ (green solid lines) are consistent with the union set $A\cup B$; local scores $c_{\lambda=1/2}(t,v)$ have similar behavior. Moreover, the scores are larger where both types of correlation are present ($A\cap B$) than where only one of the two is present ($A\setminus B$ and $B\setminus A$).
The values of the local scores $c_{\lambda}(t,v)$ with $\lambda=1$ and $\lambda=0$ are also consistent with both sets $A$ and $B$; see Figure 7. Figure 7 also shows the effect of considering $k$-hop neighborhoods with $k>1$ on local scores $c_{\lambda}(t,v)$. In uncorrelated regions, the expected value of $c_{\lambda}$ is zero. As $k$ increases, these regions become whiter (i.e., closer to zero) because aggregating over more edges reduces the score variance; see also Remark 5.7.4. We also observe a smoothing effect due to the larger overlaps between subgraphs.
AZ-whiteness test statistic: $\qquad C_{0}(\boldsymbol{\mathbf{g}}^{*})=14.7$, $\qquad C_{1/2}(\boldsymbol{\mathbf{g}}^{*})=20.8$, $\qquad C_{1}(\boldsymbol{\mathbf{g}}^{*})=14.8$. Refer to caption
AZ-whiteness test statistic: $\qquad C_{0}(\boldsymbol{\mathbf{g}}^{*})=12.0$, $\qquad C_{1/2}(\boldsymbol{\mathbf{g}}^{*})=22.4$, $\qquad C_{1}(\boldsymbol{\mathbf{g}}^{*})=14.7$. Refer to caption
Figure 9: Scores involved in the AZ-analysis of residuals on synthetic residuals from heterogeneous distributions. Residuals at different space-time locations are generated from different distributions.
Figure 10: Comparison of the proposed correlation scores with established spatial and temporal correlation measures. Top left) node scores $c_{\lambda}(v)$ and node-wise autocorrelation. Top right) time scores $c_{\lambda}(t)$ and the time-wise Moran’s II statistic. Middle) Local Moran’s II. Bottom) local scores $c_{1}(t,v)$ computed over 1-hop and 4-hop neighborhoods.